OpenStudy (anonymous):

k i need some help its a problem with f(x)-g(x) idk know how to start or what to do please help please

OpenStudy (anonymous):

Does it give you what f(x) and G(x) are

OpenStudy (anonymous):

f(x)= x^2-5 & g(x)= x+3

OpenStudy (anonymous):

ok so f(x)-g(x) x^2-5 - x+3

OpenStudy (anonymous):

so will it be -2?

OpenStudy (anonymous):

well I can't remember clearly.... if you set it equal to 0 x=-1,2

OpenStudy (anonymous):

But I don't want to give you the wrong answer I just know that first part.... sorry

OpenStudy (anonymous):

k thanks by the way .. i have no clue how to do this nothing about it

OpenStudy (anonymous):

I'm sorry... I can't remember if that is right or not

OpenStudy (anonymous):

k its ohkay

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

\[f(x) - g(x) = (x^2-5) - (x+3) = x^2-5 - x - 3 = x^2 - x - 8\]